Customers at TAB are charged for the amount of salad the take. Sampling suggests that the amount of salad taken is uniformly distributed between 5 ounces and 15 ounces. Let
Χ = Salad plate filling weight.
i. Find the probability density function of Χ
ii. What is the probability that a customer will take between 12 and 15 ounces of salad? iii. What is the probability that a customer will take fewer than 5 ounces of salad.
iv. Find Ε ( Χ) and Var ( Χ)
1
Expert's answer
2020-10-13T19:11:08-0400

solution


i) PDF of X


For a uniform distribution,


"f(x)=\frac{1}{b-a}"

For X between 5 and 15 ounces:


"f(x) =\frac{1}{15-5} =\frac{1}{10}"



Answer: "f(x)=\frac{1}{10}"


ii) Probability that "12<X<15"


"f(12<X<15)=(15-12)*\frac{1}{10} = \frac{3}{10}"



answer: "f(12<x<15)=\frac{3}{10}"

iii) probability that "X<5"


The domain of X is the range of values where the PDF is valid. The domain of "X=[5,15]"

The PDF is not valid for values of "X<5"


answer: "f(X<5) = 0"


iv) E(X) and Var(X)



"E(X) = \frac{b+a}{2}"

"= \frac{15+5}{2}= 10"




"Var(X)= \frac{(b-a)^2}{12}"

"= \frac{(15-5)^2}{12}=\frac{100}{12}=8.3333"


answer: "E(X) = 10" and "Var(X)=8.3333"